# On Discrete Morse-Bott Theory

### v1

@@ 4. Natural properties of discrete Morse-Bott functions
In this section, we observe that various properties that hold for discrete Morse functions also hold for discrete Morse–Bott functions. To state them, we demonstrate some technical lemmas as follows.
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function, $C\in\mathcal{C}_{f}$ a collection, and $\sigma\in C$ an upward noncritical cell. Then any cell $\nu\in C$ with $\nu\prec\sigma$ is upward noncritical.
@@
Since $\sigma,\nu\in C$, we have $f(\sigma)=f(\nu)$. Since $\sigma\in C$ is upward noncritical, there is a cell $\tau\notin C$ with $\tau\overset{\mathrm{nc}}{\succ}\sigma$. By definition of upward noncritical cell, we obtain $f(\sigma)>f(\tau)$. From $D^{\mathcal{C}_{f}(\tau)}(\tau)=1$ and since $\sigma\notin\mathcal{C}_{f}(\tau)$ satisfies $\tau\overset{\mathrm{nc}}{\succ}\sigma$, Lemma 2.1 implies there is a cell $\widetilde{\sigma}$ satisfying the following inequality:
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function and $C\in\mathcal{C}_{f}$ a collection. Then one of the following statements holds exclusively for any cell $\sigma\in C-C^{\mathrm{red}}$:
@@
By definition of $C^{\mathrm{red}}$, the cell $\sigma$ satisfies either (1) or (2). Assume that the cell $\sigma$ satisfies both (1) and (2). Then there are cells $\nu,\tau\notin C$ such that $\nu\overset{\mathrm{nc}}{\prec}\sigma\overset{\mathrm{nc}}{\prec}\tau$. Therefore, we have $\nu\overset{\mathrm{reg}}{\prec}\sigma\overset{\mathrm{reg}}{\prec}\tau$. By Lemma 2.1, there is a cell $\widetilde{\sigma}\neq\sigma\in\mathbb{K}$ such that $\nu\prec\widetilde{\sigma}\prec\tau$. This implies that
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function. For any cells $\nu,\tau\in\mathbb{K}$ with $\nu\overset{\mathrm{snc}}{<}\tau$, there is $\sigma\in(\nu,\tau)$ with $\nu\overset{\mathrm{snc}}{\prec}\sigma$.
@@
Put $n:=\dim\tau-\dim\nu$. We demonstrate the assertion for $\nu<\tau$ by induction on $n=\dim\tau-\dim\nu$. If $n=1$, then one can choose $\sigma:=\tau$. Thus, we may assume that $\dim\tau-\dim\nu>1$ and the assertion for any pair $\nu^{\prime}<\tau^{\prime}$ with $\dim\tau^{\prime}-\dim\nu^{\prime}<n$ holds.
By Lemma 2.2, there is a cell $\alpha\in(\nu,\tau)$. Write $m:=\dim\alpha-\dim\nu$. Suppose that $m=1$. If $f(\alpha)\leq f(\tau)$, then $f(\alpha)\leq f(\tau)<f(\nu)$ and so one can choose $\sigma:=\alpha$. Thus, we may assume that $f(\alpha)>f(\tau)$. By Lemma 2.3, there is a cell $\beta\in(\nu,\tau)$ with $\beta\nleq\alpha$. If $f(\beta)\leq f(\tau)$, then $f(\beta)\leq f(\tau)<f(\nu)$ and so the assertion holds by induction on $n$, because $\beta\in(\nu,\tau)$. On the other hand, assume that $f(\beta)>f(\tau)$. By the inductive hypothesis for $\beta<\tau$, there is a cell $\beta^{\prime}\in(\beta,\tau)$ such that $\beta\prec\beta^{\prime}$ and $f(\beta)>f(\beta^{\prime})$. Since $\beta\nleq\alpha$, we have $\beta^{\prime}\nleq\alpha$. Then $\nu<\beta^{\prime}<\tau$ and $f(\beta)>f(\tau)$. By finite iterations of this process, one can obtain a cell $\beta^{+}\in(\beta,\tau)$ with $\beta^{+}\prec\tau$, $\beta^{+}\nleq\alpha$, and $f(\beta^{+})>f(\tau)$. Then $D^{\mathcal{C}_{f}(\tau)}(\tau)>1$, which contradicts $D^{\mathcal{C}_{f}(\tau)}(\tau)\leq 1$.
Suppose that $m>1$. If $f(\alpha)\leq f(\tau)$, then $f(\nu)>f(\tau)\geq f(\alpha)$ and $\nu<\alpha$, which implies the assertion from the inductive hypothesis for $\nu<\alpha$. Thus, we may assume that $f(\alpha)>f(\tau)$. Lemma 2.2 implies that there is a cell $\widetilde{\alpha}\in(\alpha,\tau)\subset(\nu,\tau)$. The inductive hypothesis for $\nu<\widetilde{\alpha}$ implies the assertion. ∎
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function and $C\in\mathcal{C}_{f}$. Then every cell $\sigma\in\overline{C^{\mathrm{red}}}\setminus C^{\mathrm{red}}$ with $f(\sigma)=f(C)$ is upward noncritical.
@@
By $\sigma\in\overline{C^{\mathrm{red}}}$, there is a cell $\tau\in C^{\mathrm{red}}$ such that $\sigma<\tau$. Choose such a cell $\tau_{0}\in C^{\mathrm{red}}$ whose dimension is minimal. Since $f(\sigma)=f(C)$, we obtain $f(\sigma)=f(\tau_{0})$.
@@
Suppose that $\dim\tau_{0}-\dim\sigma=1$. Assume that $\sigma$ is downward noncritical. Then there is a cell $\nu\notin C$ with $\nu\overset{\mathrm{snc}}{\prec}\sigma$. By Lemma 2.3, there is a cell $\widetilde{\sigma}\in(\nu,\tau_{0})$ with $\widetilde{\sigma}\neq\sigma$. From $\tau_{0}\in C^{\mathrm{red}}$, we have $f(\widetilde{\sigma})\leq f(\tau_{0})$. By $\nu\notin C$ and $\tau_{0}\in C$, we have $f(\nu)\leq f(\widetilde{\sigma})\leq f(\tau_{0})=f(\sigma)<f(\nu)$, which is a contradiction. ∎
@@
Suppose that there is no cell $\alpha\in(\sigma,\tau_{0})$ such that $\alpha\overset{\mathrm{reg}}{\prec}\tau_{0}$. Moreover, assume that there is no cell $\beta\in(\sigma,\tau_{0})$ such that $\beta\overset{\mathrm{irr}}{<}\tau_{0}$. Then $\sigma\overset{\mathrm{reg}}{\prec}\tau$. By the finiteness of dimension, Lemma 2.2 implies the following sequence:
$\sigma\overset{\mathrm{reg}}{\prec}\beta_{1}\overset{\mathrm{reg}}{\prec}\beta_{2}\overset{\mathrm{reg}}{\prec}\cdots\overset{\mathrm{reg}}{\prec}\beta_{k}\overset{\mathrm{reg}}{\prec}\tau_{0}$
This contradicts the non-existence of $\alpha$. Thus, there is a cell $\beta\in(\sigma,\tau_{0})$ such that $\beta\overset{\mathrm{irr}}{<}\tau_{0}$.
Then $f(\sigma)=f(\tau_{0})>f(\beta)$. By $\sigma\overset{\mathrm{snc}}{<}\beta$, Lemma 4.3 implies that there is a cell $\widetilde{\sigma}\in(\sigma,\beta)$ such that $\sigma\overset{\mathrm{snc}}{<}\widetilde{\sigma}$. Therefore, $\sigma$ is upward noncritical. ∎
By the previous claim, we may assume that there is a cell $\alpha\in(\sigma,\tau_{0})$ such that $\alpha\overset{\mathrm{irr}}{\prec}\tau_{0}$.
By $\tau_{0}\in C^{\mathrm{red}}$, we obtain $f(\alpha)\leq f(\tau_{0})$. Suppose that $f(\alpha)<f(\tau_{0})$. Since $\sigma\overset{\mathrm{snc}}{<}\alpha$, by $f(\alpha)<f(\tau_{0})=f(\sigma)$, Lemma 4.3 implies that $\sigma$ is upward noncritical. Therefore, we may assume that $f(\alpha)=f(\tau_{0})$.
@@
$\sigma$ is upward noncritical.
@@
Assume that $\sigma$ is downward noncritical. Then there is a cell $\beta\prec\alpha$. By $\beta\prec\alpha\prec\tau_{0}$, Lemma 2.3 implies that there is a cell $\widetilde{\alpha}\neq\alpha$ with $\beta\prec\widetilde{\alpha}\prec\tau_{0}$. From $\tau_{0}\in C^{\mathrm{red}}$, we have the following contradiction:
@@
By the previous claim, there is a cell $\alpha\overset{\mathrm{nc}}{<}\gamma$. Since $f(\sigma)=f(\alpha)>f(\gamma)$, Lemma 4.3 implies that $\sigma$ is upward noncritical. ∎
@@
Let $f$ be a discrete Morse–Bott function. For any collection $C\in\mathcal{C}_{f}$, the set difference $\overline{C^{\mathrm{red}}}\setminus C^{\mathrm{red}}$ is a subcomplex of $\mathbb{K}$.
@@
Put $A:=\overline{C^{\mathrm{red}}}\setminus C^{\mathrm{red}}$. Fix any cells $\nu<\sigma$ with $\sigma\in A$. It suffices to show that $\nu\in A$. By definition of closure, we have $\nu\in\overline{C^{\mathrm{red}}}$. If $\nu\notin C$, then $\nu\notin C^{\mathrm{red}}$ and so $\nu\in\overline{C^{\mathrm{red}}}-C^{\mathrm{red}}=A$. Thus, we may assume that $\nu\in C$. By $\sigma\in\overline{C^{\mathrm{red}}}$, there is a cell $\tau\in C^{\mathrm{red}}$ with $\sigma<\tau$. Choose such a cell $\tau_{0}\in C^{\mathrm{red}}$ whose dimension is minimal.
Suppose that there is no cell $\alpha\in(\nu,\tau_{0})$ with $\alpha\prec\tau_{0}$. Assume that there is no cell $\beta\in(\nu,\tau_{0})$ with $\beta\overset{\mathrm{irr}}{<}\tau_{0}$. By $\sigma\in(\nu,\tau_{0})$, we have that $\sigma\overset{\mathrm{reg}}{<}\tau_{0}$. Applying Lemma 2.2 to $\sigma\overset{\mathrm{reg}}{<}\tau_{0}$, there is a cell $\sigma^{\prime}\in(\sigma,\tau_{0})$ with $\sigma^{\prime}\overset{\mathrm{reg}}{<}\tau_{0}$. By finite iterations of this process, we have a cell $\alpha\in(\nu,\tau_{0})$ with $\alpha\prec\tau_{0}$, which contradicts the non-existence of such a cell. Thus, there is a cell $\beta\in(\nu,\tau_{0})$ with $\beta\overset{\mathrm{irr}}{<}\tau_{0}$. Since $f(\beta)<f(\tau_{0})=f(\nu)$, Lemma 4.3 implies that $\nu$ is upward noncritical. By definition of $C^{\mathrm{red}}$, we obtain $\nu\in\overline{C^{\mathrm{red}}}-C^{\mathrm{red}}=A$.
Suppose that there is a cell $\alpha\in(\nu,\tau_{0})$ with $\alpha\prec\tau_{0}$. By $\tau_{0}\in C^{\mathrm{red}}$, we have $f(\alpha)\leq f(\tau_{0})$. If $f(\alpha)<f(\tau_{0})$, then Lemma 4.3 implies that $\nu$ is upward noncritical and so $\nu\in\overline{C^{\mathrm{red}}}-C^{\mathrm{red}}=A$. Thus, we may assume that $f(\alpha)=f(\tau_{0})$. By Lemma 4.4, the cell $\alpha$ is upward noncritical. Lemma 4.3 implies that $\nu$ is upward noncritical and so $\nu\in\overline{C^{\mathrm{red}}}-C^{\mathrm{red}}=A$. ∎
For any cells $\tau,\sigma\in\mathbb{K}$ and any collection $C\in\mathcal{C}_{f}$, denote by $\tau:\sigma]$ the incidence number. For any collection $C\in\mathcal{C}_{f}$, denote by $C_{k}(C^{\mathrm{red}};\mathbb{Z})$ the free $\mathbb{Z}$-module generated by the $k$-cells of $C^{\mathrm{red}}$. Put $C_{-1}(C^{\mathrm{red}};\mathbb{Z}):=\{0\}$.
@@
For any collection $C\in\mathcal{C}_{f}$ of a discrete Morse-Bott function and any integer $k\in\mathbb{Z}_{\geq 0}$, we define the boundary homomorphism
$\partial_{k}^{C}:C_{k}(C^{\mathrm{red}};\mathbb{Z})\rightarrow C_{k-1}(C^{\mathrm{red}};\mathbb{Z})$
for any cell $\tau\in C^{\mathrm{red}}$ by
$\partial_{k}^{C}(\tau):=\sum_{\begin{subarray}{c}\sigma\in C^{\mathrm{red}}\\ \sigma\prec\tau\end{subarray}}[\tau:\sigma]\sigma.$
In the previous definition, the homomorphism is well-defined, independent of the orientation of the cells.
@@
For any collection $C\in\mathcal{C}_{f}$ of a discrete Morse-Bott function and any integer $k\in\mathbb{Z}_{>0}$, the equality
@@
$X:=\overline{C^{\mathrm{red}}},\quad A:=\overline{C^{\mathrm{red}}}\setminus C^{\mathrm{red}}.$
@@
${C_{k+1}(X,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{k}(X,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{k-1}(X,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{k+1}(C^{\mathrm{red}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\partial^{C}}$${C_{k}(C^{\mathrm{red}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\partial^{C}}$${C_{k-1}(C^{\mathrm{red}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Since $X\setminus A=C^{\mathrm{red}}$, we have an isomorphism
$C_{k}(C^{\mathrm{red}})\cong C_{k}(X,A).$
@@
For any collection $C\in\mathcal{C}_{f}$ of a discrete Morse-Bott function, define the $k$-th Betti number $b_{k}^{C}$ of $C$ and the Poincaré polynomial $P_{t}(C^{\mathrm{red}})$ of $C$ as follows:
@@
We have the following characterization, which is essentially demonstrated in [knudson2015morse, p.95, Theorem 6.19].
@@
We omit the proof of the previous lemma, because it proceeds in exactly the same way as the original proof of [knudson2015morse, p.95, Theorem 6.19], except that the formula on line 19
@@
Let $f$ be a discrete Morse–Bott function. If $\nabla_{s}f$ is positively bounded, then there is a discrete Morse function $g$ with $\nabla_{s}f=-\nabla g$.

### v2

@@ 4. Natural properties of discrete Morse-Bott functions
In this section, we observe that various properties that hold for discrete Morse functions also hold for discrete Morse-Bott functions. To state them, we present the following technical lemmas.
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function and $\sigma\in\mathbb{K}$ an upward noncritical cell. Then any cell $\nu\in\mathbb{K}$ with $\nu\prec\sigma$ and $f(\nu)=f(\sigma)$ is upward noncritical.
@@
Since $\sigma\in\mathbb{K}$ is upward noncritical, there is a cell $\tau\in\mathbb{K}$ with $\tau\overset{\mathrm{snc}}{\succ}\sigma$. Therefore, we have $\nu\overset{\mathrm{reg}}{\prec}\sigma\overset{\mathrm{reg}}{\prec}\tau$. By Lemma 2.1, there is a cell $\widetilde{\sigma}\neq\sigma$ such that $\nu\prec\widetilde{\sigma}\prec\tau$. Since $D^{\mathrm{snc}}(\tau)=1$, we have $\widetilde{\sigma}\overset{\mathrm{snc}}{\not\prec}\tau$. The following inequality holds:
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function. Then one of the following statements holds exclusively for any cell $\sigma\in\mathbb{K}$:
@@
Suppose that $\sigma$ is not weakly critical. Assume that the cell $\sigma$ satisfies both (1) and (2). Then there are cells $\nu,\tau\in\mathbb{K}$ such that $\nu\overset{\mathrm{snc}}{\prec}\sigma\overset{\mathrm{snc}}{\prec}\tau$. Therefore, we have $\nu\overset{\mathrm{reg}}{\prec}\sigma\overset{\mathrm{reg}}{\prec}\tau$. By Lemma 2.1, there is a cell $\widetilde{\sigma}\neq\sigma$ such that $\nu\prec\widetilde{\sigma}\prec\tau$. This implies that
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function. For any cells $\nu,\tau\in\mathbb{K}$ with $\nu\overset{\mathrm{snc}}{<}\tau$, there is $\sigma\in(\nu,\tau]$ with $\nu\overset{\mathrm{snc}}{\prec}\sigma$.
@@
Fix any cells $\nu,\tau\in\mathbb{K}$ with $\nu\overset{\mathrm{snc}}{<}\tau$. Then $f(\tau)<f(\nu)$. Put $n:=\dim\tau-\dim\nu$. We demonstrate the assertion for $\nu\overset{\mathrm{snc}}{<}\tau$ by induction on $n=\dim\tau-\dim\nu$. If $n=1$, then one can choose $\sigma:=\tau$. Thus, we may assume that $\dim\tau-\dim\nu>1$ and the assertion for any pair $\nu^{\prime}\overset{\mathrm{snc}}{<}\tau^{\prime}$ with $\dim\tau^{\prime}-\dim\nu^{\prime}<n$ holds. By Lemma 2.2, there is a cell $\alpha\in(\nu,\tau)$. Write $m:=\dim\tau-\dim\alpha<n$.
Suppose that $m=1$. If $f(\alpha)\leq f(\tau)$, then $f(\alpha)\leq f(\tau)<f(\nu)$ and so the inductive hypothesis for $\nu\overset{\mathrm{snc}}{<}\alpha$ implies the assertion. Thus, we may assume that $f(\alpha)>f(\tau)$. Then $\alpha\overset{\mathrm{reg}}{\prec}\tau$. By Lemma 2.3, there is a cell $\beta\in(\nu,\tau)$ with $\beta\nleq\alpha$. If $f(\beta)\leq f(\tau)$, then $f(\beta)\leq f(\tau)<f(\nu)$ and so the assertion holds by induction on $n$, because $\beta\in(\nu,\tau)$. Thus, we may assume that $f(\beta)>f(\tau)$. Then $\beta\overset{\mathrm{snc}}{<}\tau$. By the inductive hypothesis for $\beta\overset{\mathrm{snc}}{<}\tau$, there is a cell $\beta^{\prime}\in(\beta,\tau]\subset(\nu,\tau]$ such that $\beta\overset{\mathrm{snc}}{\prec}\beta^{\prime}$. Suppose that $f(\beta^{\prime})\leq f(\tau)$. Then $f(\beta^{\prime})\leq f(\tau)<f(\nu)$ and so $\nu\overset{\mathrm{snc}}{<}\beta^{\prime}$. The inductive hypothesis for $\nu\overset{\mathrm{snc}}{<}\beta^{\prime}$ implies that there is $\sigma\in(\nu,\beta^{\prime}]\subset(\nu,\tau]$ with $\nu\overset{\mathrm{snc}}{\prec}\sigma$. Thus, we may assume that $f(\beta^{\prime})>f(\tau)$. Then the inductive hypothesis for $\beta^{\prime}\overset{\mathrm{snc}}{<}\tau$ implies that there is $\beta^{\prime\prime}\in(\beta^{\prime},\tau]$ with $\beta^{\prime}\overset{\mathrm{snc}}{\prec}\beta^{\prime\prime}$. Therefore, we have $\beta\overset{\mathrm{snc}}{\prec}\beta^{\prime}\overset{\mathrm{snc}}{\prec}\beta^{\prime\prime}$, which contradicts Lemma 4.2.
Suppose that $m>1$. If $f(\alpha)\leq f(\tau)$, then $f(\alpha)\leq f(\tau)<f(\nu)$ and $\nu<\alpha$, which implies the assertion from the inductive hypothesis for $\nu\overset{\mathrm{snc}}{<}\alpha$. Thus, we may assume that $f(\alpha)>f(\tau)$. Lemma 2.2 implies that there is a cell $\widetilde{\alpha}\in(\alpha,\tau)\subset(\nu,\tau)$. The inductive hypothesis on $m$ implies the assertion holds, because $\dim\tau-\dim\widetilde{\alpha}<m$. ∎
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function and $C\in\mathcal{C}_{f}$ a reduced collection. Then every cell $\sigma\in\overline{C}\setminus C$ with $f(\sigma)=f(C)$ is upward noncritical.
@@
By $\sigma\in\overline{C}$, there is a cell $\tau\in C$ such that $\sigma<\tau$. Choose such a cell $\tau_{0}\in C$ whose dimension is minimal. Since $f(\sigma)=f(C)$, we obtain $f(\sigma)=f(\tau_{0})$.
@@
Suppose that $\dim\tau_{0}-\dim\sigma=1$. Assume that $\sigma$ is downward noncritical. Then there is a cell $\nu\in\mathbb{K}$ with $\nu\overset{\mathrm{snc}}{\prec}\sigma$. By Lemma 2.3, there is a cell $\widetilde{\sigma}\in(\nu,\tau_{0})$ with $\widetilde{\sigma}\neq\sigma$. From $\tau_{0}\in C$, since $C$ is a reduced collection, we have $f(\widetilde{\sigma})\leq f(\tau_{0})$. Therefore, we have $f(\nu)\leq f(\widetilde{\sigma})\leq f(\tau_{0})=f(\sigma)<f(\nu)$, which is a contradiction. ∎
@@
Suppose that there is no cell $\alpha\in(\sigma,\tau_{0})$ such that $\alpha\overset{\mathrm{reg}}{\prec}\tau_{0}$. By $\sigma\overset{\mathrm{reg}}{<}\tau_{0}$, there is a cell $\beta\in(\sigma,\tau_{0})$ such that $\beta\overset{\mathrm{irr}}{<}\tau_{0}$. Then $f(\sigma)=f(\tau_{0})>f(\beta)$. By $\sigma\overset{\mathrm{snc}}{<}\beta$, Lemma 4.3 implies that there is a cell $\widetilde{\sigma}\in(\sigma,\beta)$ such that $\sigma\overset{\mathrm{snc}}{\prec}\widetilde{\sigma}$. Therefore, $\sigma$ is upward noncritical. ∎
By the previous claim, we may assume that there is a cell $\alpha\in(\sigma,\tau_{0})$ such that $\alpha\overset{\mathrm{reg}}{\prec}\tau_{0}$. From $\tau_{0}\in C$, we obtain $f(\alpha)\leq f(\tau_{0})$. Since $f(\tau_{0})=f(\sigma)$, if $f(\alpha)<f(\tau_{0})$, then $\sigma\overset{\mathrm{snc}}{<}\alpha$ and so Lemma 4.3 implies that $\sigma$ is upward noncritical. Thus, we may assume that $f(\alpha)=f(\tau_{0})=f(C)$.
@@
The cell $\alpha$ is upward noncritical.
@@
Assume that $\alpha$ is downward noncritical. Then there is a cell $\beta\overset{\mathrm{snc}}{\prec}\alpha$. By $\beta\prec\alpha\overset{\mathrm{reg}}{\prec}\tau_{0}$, Lemma 2.3 implies that there is a cell $\widetilde{\alpha}\neq\alpha$ with $\beta\prec\widetilde{\alpha}\prec\tau_{0}$. From $\tau_{0}\in C$, we have the following contradiction:
@@
By the previous claim, there is a cell $\gamma$ with $\alpha\overset{\mathrm{snc}}{\prec}\gamma$. Since $f(\sigma)=f(C)=f(\alpha)>f(\gamma)$, Lemma 4.3 implies that $\sigma$ is upward noncritical. ∎
@@
Let $f$ be a discrete Morse-Bott function. For any reduced collection $C\in\mathcal{C}_{f}$, the set difference $\overline{C}\setminus C$ is a subcomplex of $\mathbb{K}$.
@@
Put $A:=\overline{C}\setminus C$. Fix any cells $\nu<\sigma$ with $\sigma\in A$. It suffices to show that $\nu\in A$. By definition of closure, we have $\nu\in\overline{C}$. If $f(\nu)\neq f(C)$, then $\nu\notin C$ and so $\nu\in\overline{C}\setminus C=A$. Thus, we may assume that $f(\nu)=f(C)$. By $\sigma\in\overline{C}\setminus C$, there is a cell $\tau\in C$ with $\sigma<\tau$. Choose such a cell $\tau_{0}\in C$ whose dimension is minimal. If $f(\sigma)<f(\tau_{0})=f(\nu)$, then the fact $\nu\overset{\mathrm{snc}}{<}\sigma$ and Lemma 4.3 imply $\nu\notin C$ and so $\nu\in A$. Suppose that $f(\sigma)=f(C)$. By Lemma 4.4, the cell $\sigma\in\overline{C}\setminus C$ is upward noncritical. There is a cell $\sigma^{\prime}\in\mathbb{K}$ with $\sigma\prec\sigma^{\prime}$ and $f(\sigma)>f(\sigma^{\prime})$. Since $f(\sigma^{\prime})<f(\sigma)=f(C)=f(\nu)$ and $\nu<\sigma\prec\sigma^{\prime}$, we have $\nu\overset{\mathrm{snc}}{<}\sigma^{\prime}$. Lemma 4.3 imply $\nu\notin C$ and so $\nu\in A$. Thus, we may assume that $f(\sigma)>f(C)$. Then $f(\sigma)>f(C)=f(\tau_{0})=f(\nu)$.
Claim 4.
There is a cell $\alpha\in(\sigma,\tau_{0})$ such that $f(\alpha)\leq f(\tau_{0})$ and $\alpha\notin C$.
Proof.
Assume that there is no cell $\alpha\in(\sigma,\tau_{0})$ such that $f(\alpha)\leq f(\tau_{0})$ and $\alpha\notin C$. By $\tau_{0}\in C$, we have that $\sigma\overset{\mathrm{reg}}{<}\tau_{0}$ and $\dim\tau_{0}-\dim\sigma\geq 2$. Applying Lemma 2.2 to $\sigma\overset{\mathrm{reg}}{<}\tau_{0}$, there is a cell $\alpha_{0}\in(\sigma,\tau_{0})$. By $\tau_{0}\in C$ and $\alpha_{0}<\tau_{0}$, we obtain that $\alpha_{0}\in A$, because $\tau_{0}>\sigma$ is a minimal cell which is in $C$. By the assumption, we obtain that $f(\alpha_{0})>f(\tau_{0})$. Then $\alpha_{0}\overset{\mathrm{snc}}{<}\tau_{0}$. Lemma 4.3 imply that there is a cell $\alpha_{1}\in(\alpha_{0},\tau_{0})\subset(\sigma,\tau_{0})$ with $\alpha_{0}\overset{\mathrm{snc}}{<}\alpha_{1}$. By the assumption and the minimality of $\tau_{0}$, we obtain that $\alpha_{1}\overset{\mathrm{snc}}{<}\tau_{0}$. By finite iterations of this process, one can obtain a cell $\alpha^{+}\in(\sigma,\tau_{0})$ with $\alpha^{+}\overset{\mathrm{snc}}{\prec}\tau_{0}$, which contradicts that $\tau_{0}$ is weakly critical. ∎
If $f(\alpha)<f(\tau_{0})$, then the fact $\nu\overset{\mathrm{snc}}{<}\alpha$ and Lemma 4.3 imply $\nu\notin C$ and so $\nu\in A$. Thus, we may assume that $f(\alpha)=f(\tau_{0})$. Then $f(\alpha)=f(\tau_{0})=f(C)=f(\nu)$. By $\tau_{0}\in C$ and $\alpha<\tau_{0}$, we obtain that $\alpha\in A$, because $\tau_{0}>\sigma$ is a minimal cell which is in $C$. Since $f(\alpha)=f(C)$ and $\alpha\in A=\overline{C}\setminus C$, Lemma 4.4 implies that $\alpha\in\overline{C}\setminus C$ is upward noncritical. There is a cell $\alpha^{\prime}\in\mathbb{K}$ with $\alpha\prec\alpha^{\prime}$ and $f(\alpha)>f(\alpha^{\prime})$. Since $f(\alpha^{\prime})<f(\alpha)=f(C)=f(\nu)$ and $\nu<\alpha\prec\alpha^{\prime}$, we have $\nu\overset{\mathrm{snc}}{<}\alpha^{\prime}$. Lemma 4.3 imply $\nu\notin C$ and so $\nu\in A$. ∎
For any cells $\tau,\sigma\in\mathbb{K}$, denote by $\tau:\sigma]$ the incidence number. For any reduced collection $C$ and any integer $k\in\mathbb{Z}_{\geq 0}$, denote by $C_{k}(C;\mathbb{Z})$ the free $\mathbb{Z}$-module generated by the $k$-cells of $C$. Put $C_{-1}(C;\mathbb{Z}):=\{0\}$.
@@
For any reduced collection $C\in\mathcal{C}_{f}$ of a discrete Morse-Bott function $f$ and any integer $k\in\mathbb{Z}_{\geq 0}$, we define the boundary homomorphism
$\partial_{k}^{C}:C_{k}(C;\mathbb{Z})\rightarrow C_{k-1}(C;\mathbb{Z})$
for any $k$-cell $\tau\in C$ by
$\partial_{k}^{C}(\tau):=\sum_{\begin{subarray}{c}\sigma\in C\\ \sigma\prec\tau\end{subarray}}[\tau:\sigma]\sigma.$
In the previous definition, the homomorphism is well-defined and independent of the orientation of the cells.
@@
For any reduced collection $C\in\mathcal{C}_{f}$ of a discrete Morse-Bott function $f$ and any integer $k\in\mathbb{Z}_{>0}$, the equality
@@
$X:=\overline{C},\quad A:=\overline{C}\setminus C.$
@@
${C_{k+1}(X,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{k}(X,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{k-1}(X,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{k+1}(C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\partial^{C}}$${C_{k}(C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\partial^{C}}$${C_{k-1}(C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Since $X\setminus A=C$, we have an isomorphism
$C_{k}(C)\cong C_{k}(X,A).$
@@
For any reduced collection $C\in\mathcal{C}_{f}$ of a discrete Morse-Bott function $f$, define the $k$-th Betti number $b_{k}^{C}$ of $C$ and the Poincaré polynomial $P_{t}(C)$ of $C$ as follows:
@@
We have the following characterization, which is essentially demonstrated in [9, p.95, Theorem 6.19].
@@
We omit the proof of the previous lemma, because it proceeds in exactly the same way as the original proof of [9, p.95, Theorem 6.19], except that the formula on line 19
@@
Let $f$ be a discrete Morse-Bott function. If $\nabla_{s}f$ is positively bounded, then there is a discrete Morse function $g$ with $\nabla_{s}f=-\nabla g$.